OpenStudy (anonymous):

Solve the system x + y = 6 and x - y = 2. What is the value of x? Answer 6 2 4 0

OpenStudy (anonymous):

solve for y on the first equation and then substitute it in the second equation, will leave you with x's and then just solve for x

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

ok lets start with x + y = 6 we begin by moving the x to the right so lets subtract -x to both sides x + y - x = 6 - x => y = 6 - x now lets use that to substitute for the y on the second so

OpenStudy (anonymous):

so x - ( 6 - x ) = 2 => x - 6 + x = 2 => 2x - 6 = 2 2x - 6 + 6 = 2 + 6 => 2x = 8 => x = 4

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

I did it step by step so you can follow how I did it but the answer is x = 4

OpenStudy (anonymous):

alright thank you so much!